If 15 x's = 20 y's and 16 y's = 10 z's, then 6 x's is how many z's?
Answer by stanbon(75887)

If 15 x's = 20 y's and 16 y's = 10 z's, then 6 x's is how many z's?
-----------------------
x = (4/3)y
z = (8/5)y
-------------
Let y = 15
Then x = 20
And z = 24
----------------
EQUATION:
6x = 6^20 = 120
120/z = 120/24 = 5
--------------------
6x = 5z's
===============
Cheers,
Stan H.